OpenStudy (anonymous):

Solve –3x2 + 30x – 90 = 0

OpenStudy (anonymous):

–3x2 + 30x – 90 = 0 -3(x^2-10x+9)=0 or x^2-10x+9=0 x^2-9x-x+9=0 x(x-9)-1(x-9)=0 (x-9)(x-1)=0 or x-9=0 or x-1=0 hence x=9 or x=1